OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A: f(x) = 25(1.08)x B: x f(x) 0 22.00 1 21.56 2 21.13 C: The initial price of this stock is $30, but it has been increasing 4% each day. 1.) Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. 2.) Graph the price function for Stock D. Explain the key features of the graph wit

OpenStudy (anonymous):

NEED HELP!!!!! MEDALS!!!!! WILL BE YOUR FAN!!!!!

OpenStudy (anonymous):

@MERTICH #1 is h(x)=30+(0.04×30)x=30+1.2x right?

OpenStudy (anonymous):

I messed up, gonna make a new one, ok??

undeadknight26 (undeadknight26):

do u still need help here?

OpenStudy (anonymous):

Yes

undeadknight26 (undeadknight26):

ok one sec im afk.

OpenStudy (anonymous):

I am working on it again. I was struggling to be online. Am re-looking at it again. TODAY YOU WILL GET YOUR ANSWER, DON'T WORRY!

undeadknight26 (undeadknight26):

For this question you have shown you have completed Part 1 right? If so all you have to do from there is create a function for c.

undeadknight26 (undeadknight26):

since it started at 30 your first step will be to create the first part of ur exponential function: f(x) = 30(???) since we know it increases by 4%. so f(x) = 30(1.4)x